Meeting 22-4-2015¶
Date: 22nd April 2015
Time: 15:00 UK time
Attendance: Chris Heidt, Durga, Melissa, Ed, Adam, Pierrick, Craig, Paul, Ray
Apologies:
Agenda and Minutes¶
Actions from Last Meeting¶
- Look at docs to determin what to release as MICE notes and how - MU - ONGOING
- Kalman doesnt work on straights. - Chris and Adam looking to fix _ Adam says that Chris hunt has made solid progress but it is not finished yet. He has gone back to the beginning and is modifying in an easy and reproducible way.
- Not yet able to test against multiple tracks per trigger with MC - Durga - Melissa to remind Durga - DONE TBC June
- Real data unpacking broken - Paul -> Test data taken the other week does unpack StepIV data unpacker but it does not have data in. Adam has manged to get it to reconstruct digits but nothing more and it is very slow. Ed has updated the issue tracker with some hints (he thinks that everything is somehow being stored as digits) but there is still a lot to do and quite a few unknowns. We must have an update in two weeks on a clear description of the issue.
- Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick - Ongoing
- Add to John Govans' rota for daily walk round to test our setup. - Pierrick - Ongoing
- write emergency shutdown procedure - MU please review at http://micewww.pp.rl.ac.uk/projects/tracker/wiki/TrackerOperationsDocs
1. Status of hardware, commissioning work and ongoing plan EO, AB, CMW, PR, MU¶
*Alan has been here since last Wednesday and Paul R on Thursday that week. Ed with them began by looking at the noise.*Noise issue is now fixed by reducing the size of the integration window (below 175ns wide).
- the down side is that we are integrating charge for less time which effects how much of the beam we can take. This is being investigated.
- Calibration was next, and all cryostats were done. Going well.
- James's code is merged with IOC and is now working well. Into that system low level socket access exists so that Paul R can communicate with front end boards.
- Paul updated the spreadsheet to use the socket above. Spreadsheet can now validate and update firmware on the boards.
- Friday night it was discovered that outdated firmware was used on some spare boards. This is now facilitated by above point.
- Temperature regulation and readout now working across the board.
- Alan is currently extending the length of the lemo cables so that they are not so stretched.
- Ed and Yordan have decided some quite significant changes (wrapper reducing) of Davids code to make it more useable. Timescale of a few weeks.
- bias calibration needs to be done on all four cryostats. It appears to be working but we are not sure if it is believable. Need to go round all 4 with a pulser box. This will happen on next visit (1 day per cassette so approx 1 week).
- Ed will look at DAQ code while in Sheffield (1 week or so)
- On cryo 3 one temp resistor has gone bad. Paul R says its ok, as the boards are designed for this and we can regulate off the neighbouring sensor. However this functionality isn't in the code and must be added.
- Integration with tracker still to get working, it was tested on single station but not recently, inc cabling DATE code reading out, alive window in the trigger, validate trigger firmware etc (approx 2 weeks).
- After waveguides etc we are ready to set up with beam (approx 1 week).
- cryo 3 has a very small leak on it as pressure is dropping - Craig will sort.
- Vacuum pump failed on Monday.
- Have a spare Turbo pump but pare vacuum pump used. Crag will order a full suite of spares.
- cryo2 wont quite get cool enough and there is a temperature differential between the two cassettes. The idea is to cool it more to solve this. It has never been below 6.5 and needs to be 6.2. This will be done by increasing rate to 60Hz rather than 50Hz, for this an AC inverter may be needed. Around 100 pounds and a few days to put together.
- Cryo lids still needed. Melissa to ask Geoff or Kevin.
*Interlocks: DL and Craig will work on it next week (with Ian and John Webb). One of the crates will have to go to DL. About two weeks to near completion inc implementing the VESDAs.
2. Software: PK, AD, CH¶
In particular this week:- Data unpacking - next week see actions
- Kalman - next week see actions
- DAQ in data structure - see above
- CDB
Geometry, Calibrations, Configurations, CDB - C. Heidt¶
- Meeting to happen tomorrow add cabling says Durga.
Monte Carlo¶
- TALK - Update on MC from Chris Heidt - 20mins
- Notes:
- Fibres are only implemented to the edge of the stations they do not wrap around.
Pattern Recognition - AD,¶
No update
Online Displays - MU¶
*3/4 of the plots are now in and working and were used for mock data challenge. The other are the non-immediately-essentials and are coming soon.
Performance and Analysis -PK¶
- Slow running - Difficult to understand as it is compounded with unpacking. Yordan has managed to run the code using his framework and there doesn't seem to be speed problems then. It is on the list but the list is long.
- MAUS - Adam - removing strings and other works to speed up MAUS in general. String conversions happen every time data is passed between maps and wrappers.
- These issues cannot be unravelled until after unpacking.
Issues¶
As discussed.
Tracker Software Paper - AD, PK DR¶
Ongoing but there is an updated later draft which has been heavily updated and trimmed by Paul K.
Documentation - PK, AD, C. Heidt¶
- Ed C and M documentation needed. See actions from ealier.
3. What and when are there opportunities to do before step IV and what can we do to be ready for them?¶
- Next ISIS run. Are we able to go ahead with the last ISIS weekend run?
- To synchronise with the ISIS RF which was done at the single station test in two days (but we can try to complete in 1) and we will need a day with LEDs in advance. One station in one tracker at least necessary for this but one in each trackers is best.
- And to ensure that everything not completed as part of the mock data challenge is completed.
4. Optimising the DAQ and Controls and Monitoring - EO, AB, PR¶
- C and M has been sped up making the board config 4x faster.
- It also now can only configures things are not configured but this must be added to the gui.
- an alarm if config gets lost should be added as well as the ability to turn on/off heaters on vlpcs.
5. MICE hall, inc Services hardware and requirements : CM okay¶
Helium line coming this week, which will mean we wont have to keep swapping bottles.
6. Step IV Analysis Readiness - MU¶
7. Risks¶
7.1 Hardware¶
- Daresbury electronics: vacuum, tracker interlocks
- Reforming external waveguides (could lead to breakage)
- Northside PRY (not us, but knock effect to us)
7.2 Software¶
- Not yet able to test against multiple tracks per trigger with MC -Action for Durga
- Real data unpacking broken DA
8. Tracker Documentation¶
We need to produce documentation for the control room and ISIS, for Steve B. We cannot necessarily get that far with operations and startup as we are not yet working with the CandM as we will but we can write something and finalise safety details.
8.1 Startup Procedure¶
8.2 Operations Procedure¶
8.3 Tracker Online Emergency Procedure¶
- We have been asked to provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
Documentation to Include:
- schematics of the system
- a list of associated checklists at power on, and who works through them.
- a list of instructions to shifters on how to operate and monitor the system
- emergency shutdown instructions
- power-up instructions
- trouble-shooting : what to watch for
- a list of regular maintenance and safety checks
- A beta version is available at http://micewww.pp.rl.ac.uk/projects/tracker/wiki/TrackerOperationsDocs - MU
h3.
5. Other Items
- Installation Manual - GB MU - ongoing
- LED system - EO - External used for calib, Internal not used since Dec but will be used after gain and pedestals determined.
- Hall probes - MU - Ongoing
AoB¶
*Software expert weekly meetings discussion - Paul K - Passed: 15 minutes on the off weeks from Tracker meetings.
New Actions¶
- Kalman doesnt work on straights. - Chris and Adam looking to fix
- Real data unpacking broken. A proper summary of the issue, its status and timeline to be presented next week. - Paul.
- Software update from Paul next meeting - PK.
- Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick
- Add to John Govans' rota for daily walk round to test our setup. - Pierrick
- Look at docs to determin what to release as MICE notes and how - MU - ONGOING
- Ed to talk to Chris Heidt about CDB so that he can determin method needs, cabling, etc Pierrick and Durga to come also. Chris will then Meet with Janusz later in the week.
- Update in depth Traker schedule and agree with Ed and Craig. - MU
- Cryo lids still needed. Melissa to ask Geoff or Kevin
- Add his version of the code that can look at single station test data to launchpad branch with an appropriate name - Chris Heidt
- Put the above on the data section of the Tracker wiki. -MU
Phone Details¶
- The four digit company id: 4673 (callers from outside the UK)
- The pin number: 419357
Phone numbers:
UK (Imperial College --preferred) | 0845 545 4460 | English |
Austria | 0820 4000 1503 | English |
0820 4000 15 02 | German | |
Belgium | 070 35 99 45 | English |
070 35 99 94 | Dutch | |
070 35 98 66 | French | |
Canada | 1 712 432 2832 | English |
1 712 432 2844 | French | |
Germany | 01803 001 178 | English |
01570 248 8303 | English (mobile) | |
01803 001 177 | German | |
01570 248 8101 | German (mobile) | |
Ireland | 0818 270 007 | English |
Italy | 848 390 166 | English |
848 391 819 | Italian | |
Netherlands | 0870 001 901 | English |
0870 001 909 | Dutch | |
Poland | 0801 003 543 | English |
0801 003 533 | Polish | |
South Africa | 087 550 0375 | English |
Spain | 902 885 318 | English |
902 881 200 | Spanish | |
Sweden | 0939 2066 400 | English |
0939 2066 300 | Swedish | |
Switzerland | 0848 560 190 | English |
0848 560 378 | French | |
0848 560 347 | German | |
0848 560 385 | Italian | |
United Kingdom (PowWowNow) | 0844 4 73 73 73 | English |
United States | 1 415 363 0833 | English |
Updated by George, Melissa over 8 years ago ยท 13 revisions